[ A long wire is cut into three smaller pieces in the ratio of 7:3:2. if the shortest piece is 16 cm, what is the area of the rectangle with the largest area that can be created using the longest piece? ]
Answer by TimothyLamb(4379)

a = (7/12)L
b = (3/12)L
c = (2/12)L
---
c = 16 cm
c = (2/12)L
16 = (2/12)L
---
L = 16(12/2)
L = 96 cm
---
a = (7/12)L
a = (7/12)96
a = 56 cm
---
b = (3/12)L
b = (3/12)96
b = 24 cm
---
a is the longest piece at 56 cm
---
what is the area of the rectangle with the largest area that can be created using the longest piece?
it's actually a square, so:
L = w = (1/4)56
L = w = 14
a = Lw = 14^2
a = 196 sq.cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php